Given $7m and the following scenario: 180-day US interest rate = 4.5% 180-day Japanese interest rate = 3% Current Spot rate = 115.75 (¥/$) 180-day Forward rate = 112.50 (¥/$) The covered interest arbitrage profit potential is: a. 5.78% b. 2.0% c. 4.28% d. 1.5% e. None of the Above
Answer #1

An arbitrageur executes a covered interest arbitrage strategy by exchanging domestic currency for foreign currency at the current spot exchange rate, then investing the foreign currency at the foreign interest rate.

We have $7M

we invest it in USA so the deposit rate is 4.5% 180 day (ignoring compounding effect)

$7 * 4.5% * 180/365 = .16 M$

or we can exchange the $ with yen at current spot rate we get

$7M * 115.75 = 810.25 M yen

Now we invest it in japan where interest rate is 3% 180 days

810.25 * 3% * 180 / 365 = 11.99 m yen

Future value = 822.24 M yen

180 day forward rate = 112.5 yen/ $

so the Future value in US $ = 822.24/112.5 = 7.31 M $

so the profit is in exchange the $ into yen and invest in japan convert it into forward rate =.31

Interest arbitrage profit = .31/7 = 4.28%(approx)

Covered interst arbitrage profit = C 4.28%